# UrbanKGent: A Unified Large Language Model Agent Framework for Urban Knowledge Graph Construction

### v1

@@ 5.1.4. Evaluation Protocol
${}^{\spadesuit}$ 0.38/4.17 0.42/3.98 0.37 0.34
${}^{\star}$ 0.34/4.06 0.45/4.02 0.34 0.39
${}^{\ddagger}$ 0.45/4.32 0.40/3.97 0.45 0.23
${}^{\dagger}$ 0.44/4.10 0.47/3.85 0.46 0.43
5.2. Main Result (RQ1)
The performance results are reported in Table 3. As can be seen, the constructed agent outperforms all twenty-one baseline models on two UrbanKGC datasets. Specifically, the UrbanKGent-13B achieves (15.56%, 14.29%, 14.89%, and 11.90%) improvements compared with the state-of-the-art GPT-4 with the same inference pipeline on NYC. The improvements on CHI are (15.22%, 17.07%, 13.46%, and 13.95%), respectively. Moreover, the UrbanKGent-7B also achieves comparable performance compared with the GPT-4. Such improvement demonstrates the superiority of the constructed UrbanKGC agent.
Meanwhile, we observe that the zero-shot LLMs perform poorly in the UrbanKGC tasks, even using GPT-4. In addition, although the demonstrations provided by In-context-learning can incorporate the UrbanKGC task information, the performance gain is limited and even leads to performance degradation in smaller LLMs (e.g., Llama-2-7/13B). Besides, we find that fine-tuning LLMs can make obvious improvements in the overall performance. Through vanilla fine-tuning, the Llama-2-7B and Llama-2-13B could achieve comparable performance with GPT-3.5 under the zero-shot reasoning and In-context learning settings.
Moreover, although the LLMs (i.e., Llama-2-7B and Llama-2-13B) using the UrbanKGent inference pipeline perform slightly worse than the vanilla fine-tuning method, they could obtain better performance compared with zero-shot reasoning and In-context learning paradigms. Such results demonstrate the benefit of knowledgeable instruction design and external tool innovation, but also indicate its performance bottleneck (i.e., well-designed UrbanKGent prompting pipeline unable to beat vanilla fine-tuning). As a deeper exploration, our work fills this gap through hybrid instruction fine-tuning, and the fine-tuned UrbanKGC agents, whether 7B or 13B, can achieve state-of-the-art performance in UrbanKGC tasks.
Table 5. Comparison among LLM-based UrbanKGC methods in four ways.
@@ 5.2. Main Result (RQ1)
5.3. In-Depth Analysis (RQ2)
We conduct an in-depth analysis of the proposed instruction generation and tool-augmented iterative trajectory refinement module on the NYC dataset. Specifically, for the RTE and KGC task, we validate the effectiveness of each block by comparing the following variants: (1) ${}^{\spadesuit}$ removes knowledgeable instruction template in RTE and KGC task; (2) ${}^{\star}$ removes multi-view design in RTE task; (3) ${}^{\ddagger}$ removes tool invocation; (4) ${}^{\dagger}$ removes iterative trajectory self-refinement. We summarize the results in Table 4, and obtain the following observations.
Figure 5. The model latency and cost of constructed UrbanKGent-13B and GPT-4 in UrbanKGC tasks. We report the total inference time and cost of 1,000 RTE and KGC tasks.
5.4. Complexity and Effiency Analysis
@@
Moreover, we also provide efficiency analysis in Figure 5. As can be seen, UrbanKGent-13B achieve lower inference speed in latency and reduce the cost by roughly 20 times in both RTE and KGC tasks. The detail is in Appendix H.
5.5. Agent Application and Deployment (RQ3)
We have deployed a prototype system [^10] https://htmlpreview.github.io/?https://raw.githubusercontent.com/usail-hkust/UrbanKGent/main/UrbanKGent%20Demo/index.html to facilitate real-world UrbanKGC construction. The service is equipped with the UrbanKGent-13B version and is currently optimized for New York City and Chicago.
As shown in Table 6, compared to the previous UrbanKG benchmark (Ning et al., 2023), we only use roughly one-fifth of the data for constructing the UrbanKGs with the same scale of triplets and entities, and even expanding the variety of relationships to a thousand times the original types. Please refer to Appendix F for more details.
Table 6. Statistics comparision between constructed two UrbanKGs in NYC and CHI and existing UrbanKG dataset and benchmark UUKG (Ning et al., 2023). For relation and data volume, we report the increase and decrease percentages in terms of absolute percentage returns.
Dataset # Entity # Relation # Triplet Data Volume
NYC-Large 265,938 24,052 ($\uparrow 1850\%$) 975,102 39,600 ($\downarrow 597\%$)
CHI-Large 191,473 21,089 ($\uparrow 1622\%$) 610,258 28,378 ($\downarrow 495\%$)
NYC-UUKG 236,287 13 930,240 236,277
CHI-UUKG 140,602 13 564,400 140,577
6. Related work
Our work is related to domain-oriented agent construction, LLMs for knowledge graph construction and urban knowledge graph. We briefly discuss them in this section
6.1. Domain-Oriented Agent Construction
The concept of language agent (Chen et al., 2023) has become very popular recently, and a variety of LLM agents targeting different domains have been proposed. For example, Voyager (Wang et al., 2023c) is constructed for automated game exploration, WebGPT (Gur et al., 2023) is an HTML agent for diverse document understanding tasks, LLMLight (Lai et al., 2023) constructs a language agent for transportation domain, K2 (Deng et al., 2023), GeoGalactica (Lin et al., 2023) and GeoLLM (Li et al., 2023c) propose to re-train language agent for geospatial semantic understanding. In addition, many recent works like Auto-GPT (aut, 2023) and CAMEL (Li et al., 2023a) aim at proposing an autonomous agent framework for agent construction. Nevertheless, there is still no UrbanKGC agent construction framework for the urban computing domain.
6.2. LLMs for Knowledge Graph Construction
Knowledge Graph Construction aims to extract structural information from unstructured text (Ye et al., 2022). Recently, the advent of Large Language Models (LLMs) invigorated the field of Natural Language Processing (NLP). Many studies have begun to explore the potential of LLMs in the domain of knowledge graph construction. For example, (Wei et al., 2023; Li et al., 2023b) finds that transforming the NER and RE task into a multi-turn question-answering dialog could improve the performance of LLMs on the KG construction task. (Xie et al., 2023) explicitly derive syntactic knowledge to guide LLMs to think, which could develop the performance of NER. Despite these LLM-driven knowledge graph construction methods (Yuan et al., 2023; Lu et al., 2023) in general domains being widely investigated, knowledge graph construction in specific domains still remains an open challenge (Zeng and Wang, 2023). There is still no promising knowledge graph construction method in the urban computing domain, and this hinders its potential achievement.
6.3. Urban Knowledge Graph
Urban knowledge graph has been proven useful in various urban tasks, such as traffic flow prediction (Yang and Qi, 2022; Tan et al., 2021; Liu et al., 2021; Zhao et al., 2020), mobility prediction (Wang et al., 2021), site selection (Liu et al., 2022), city profiling (Zhou et al., 2023), crime prediction and so on (Ning et al., 2023). Their common approach involves manually extracting urban entities and defining urban relations to construct an urban knowledge graph. For example, (Wang et al., 2021) construct a dedicated spatiotemporal knowledge graph regarding trajectory and timestamp as entities to improve trajectory prediction and (Liu et al., 2022) construct user check-in relations to help mobility prediction. Nevertheless, existing UrbanKGs heavily rely on manual design, leading to high labor costs and constrained by a limited variety of urban entities and relation categories.
7. Conclusion
In this work, we proposed UrbanKGent, the first automatic UrbanKG construction agent framework with Large Language Models (LLMs). We first constructed a knowledgeable instruction set to adopt LLMs for different UrbanKGC tasks. Then, we proposed a tool-augmented iterative trajectory refinement module to facilitate the instruction tuning of various large language models. Extensive experimental results demonstrate the advancement of UrbanKGent in improving UrbanKGC tasks. We release the UrbanKGent-13B, an agent based on Llama-13B, with lower latency and cost compared with GPT-4 for UrbanKG construction. We hope the open-source UrbanKGent can foster future urban knowledge graph research and broader smart city applications.
8. Limitation and Future Work
This work has limitation on the further application demonstration of construction UrbanKGs, although proposed UrbanKGent-13B could construct a UrbanKG with a thousandfold relationship using only one-fifth of data. In addition, the evaluation method in this work is cost-intensive although GPT evaluation and Human evaluation has been experimentally demonstrated to be consistent. Despite the above limitations, we hope the opensource UrbanKGC agent can foster more extensive UrbanKG research and broad smart city application. In the future, we will derive extra image-modality data to further enrich UrbanKGC.
Appendix A UrbanKGC Construction
Table 7. The detailed statistic of RTE datasets. We report the maximum length, minimum length, and average length of urban text in the RTE dataset.
Dataset Max Length Min Length Avg Length # Records
NYC-Instruct 1,747 68 437 232
CHI-Instruct 1,120 25 408 122
NYC 2,708 51 433 2,089
CHI 1,883 32 445 1,102
NYC-Large 4,277 28 1,167 39,600
CHI-Large 4,306 36 804 28,378
This section presents detailed UrbanKGC dataset statistic information for relational triplet extraction (RTE) and knowledge graph completion (KGC) tasks.
To facilitate the understanding of the constructed RTE dataset, we summarize the distribution of entity types in the four RTE datasets, and the text statistics of entity textual descriptions. As shown in Table 7, the entity distribution and textual statistics of the instruct dataset and test data are similar. In addition, the record in dataset (i.e., NYC-Instruct and CHI-Instruct) used for instruction tuning is not overlapping with that in test dataset (i.e., NYC and CHI). This can avoid potential data leakage issues.
As for the KGC dataset, we provide illustrative visualization of the five RCC relationships (Mani et al., 2010) in Figure 6 for better understanding. Specifically, the disconnected (DC), externally connected (EC), partially overlapping (PO), equal (EQ), tangential and non-tangential proper parts (IN) together depict the basic geospatial relationship between urban entities. Moreover, to facilitate the understanding of the constructed KGC dataset, we visualize the geometry range of the head entity and the tail entity in four KGC datasets. As can be seen in Figure 7, the entity distribution in the instruct dataset (i.e., NYC-Instruct and CHI-Instruct) and test dataset (i.e., NYC and CHI) are similar. Due to the KGC task follows a zero-shot setting, we are unable to provide accurate distribution proportions of the 5 RCC relationships in the four datasets. However, it is intuitive can be seen that all four datasets contain the five RCC relationships shown in Figure 6, which provides a guarantee for the practical significance of KGC task in this work.
Figure 6. Given two geo-entities X and Y, the illustrative visualization of five types of RCC relationships. In this work, we consider entities with coordinate geometry to a circles with very small radii.
Appendix B Instruction Template
B.1. Instruction Template in UrbanKGent
Figure 7. The geometry range visualization of the head entity and tail entity of four KGC datasets. The horizontal and vertical coordinates are longitude and latitude, respectively. The blue and red polygons stand for entities with the polygon geometry, the purple line string stands for the entities with linestring geometry and the green point is for the coordinate entities.
B.2. Instruction Template in Baselines
LLMs-based In-context Learning Methods. We first construct several few-shot demonstrations via the chain-of-thought prompting techniques (Wei et al., 2022), which is popular for automatic demonstration generation. The detailed instruction template is shown in Figure 9. Then, we add these demonstrations before the test question, and the detailed instruction template is shown in Figure 9.
Figure 8. The reasoning process of UrbanKGent-13B and the detailed instruction template for UrbanKent inference pipeline. The content in the gray box is the instruction, and the content in the green box is the agent’s response.
Figure 9. The instruction template of zero-shot reasoning, In-context learning and vanilla fine-tuning baselines.
Table 8. The detailed geospatial tool name, description, interface input, and output. Each tool interface is implemented by python and it is self-programmed by GPT-4.
Tool name Tool description Input Output
Geohash Geohash encoding Geometry Geohash code (8-bit)
Appendix C Geospatial Toolkit
Table 9. Illustrative RTE evaluation example when we utilize human evaluation and GPT evaluation method. We calculate the accuracy by counting the proportion of true triplets. The label for GPT evaluation method is invisible.
Type Urban text Results Label
Number of the
true triplet
Number of the
false triplet
Human
Evaluation
<Columbia University,
locate-in, New York City >
1 0
GPT
Evaluation
- 1 0
Appendix D Evaluation
This section presents the detailed evaluation process and examples to help the reader have a better understanding. Moreover, we provide a comprehensive analysis to demonstrate why GPT and Human evaluations are highly aligned.
D.1. Human Evaluation Process
For the relational triplet extraction (RTE) task, we provide an evaluation example in Table 9. Given the urban text sentence ”Columbia University is a private Ivy league research university in New York City.”, the human annotators are required to first label the triplet contained in this urban text, described as triplet ¡Columbia University, locate-in, New York City¿. Based on the label, then, they are instructed to evaluate how many true triplets and false triplets in the results from the models. Finally, they will fill out the evaluation form (i.e., the number of the true triplets and the number of the false triplets). We will calculate the accuracy of results based on these annotated forms.
For the knowledge graph completion (KGC) task, given the head entity ¡Columbia University, Point (40.8075, 73.9626)¿ and the tail entity ¡Empire State Building, Point (40.7484, 73.9857)¿, the human annotators are first required to complete their missing geospatial relationship from the five relation candidate (i.e., DC, EC, PO, EQ and IN). Specifically, the annotation could be achieved by manually visualizing the location of two entities given on the map and following the RCC relation rule in Figure 6 to determine their geographical relationships. Finally, they will fill out the evaluation results (i.e., True/False). We will calculate the accuracy based on these evaluation results.
D.2. GPT Evaluation Process
For the relational triplet extraction (RTE) task, we provide an evaluation example in Table 9. Given the urban text sentence ”Columbia University is a private Ivy league research university in New York City.”, and the model extraction results (i.e., ¡Columbia University, locate-in, New York City¿), we directly prompt GPT-4 to fill out the evaluation form (i.e., the number of the true triplets and the number of the false triplets). Then, the accuracy could be obtained based on these self-evaluated results.
For the knowledge graph completion (KGC) task, given the head entity ¡Columbia University, Point (40.8075, 73.9626)¿ and the tail entity ¡Empire State Building, Point (40.7484, 73.9857)¿ and completed geospatial relationship, we directly prompt GPT-4 the justify if the results are correct. Specifically, we will first explicitly call eight external geospatial tools in Table 8, and combine all these eight calculation results into the prompt. Then we feed these prompts into GPT-4 to help it make a comparable accurate evaluation. Finally, we obtain the evaluation results (i.e., True/False) for every record. We will calculate the accuracy based on these evaluation results.
D.3. Evaluation Consistency
We conduct experiments to assess the consistency between GPT-4’s evaluation results and human evaluation results. In this work, we follow GPTScore (Fu et al., 2023a) to use Spearman coefficient (Zar, 2005) to investigate the correlation between GPT evaluation and human evaluation. Based on the evaluation results in Tabel 3, we can conduct correlation analysis separately by the types of model and paradigm.
Table 10. The average Spearman correlation value between human evaluations and GPT-4 evaluations. ”Repeat X times” refers to instructing GPT-4 to generate judgments X times, and adopting the answer that appears most frequently (e.g., True/False for the KGC task and Number of the true triplet for the RTE task) as the final decision.
Task Repeat Once Repeat 3 times Repeat 5 times Repeat 10 times
RTE 98.56% 87.31% 86.42% 85.27%
KGC 94.86% 91.23% 90.09% 91.24%
Appendix E Case Study
@@
Figure 10. The Spearman correlation between the GPT evaluation and human’s evaluation under five different LLM backbones (i.e., Llama-2-7B, Llama-2-13B, Llama-2-70B, GPT-3.5 and GPT-4).
Figure 11. The Spearman correlation between the GPT evaluation and human’s evaluation under four LLM paradigms (i.e., Zero-shot learning (ZSL), In-context learning (ICL), Vanilla fine-tuning (VFL) and UrbanKGent inference).
Appendix F Prototype System
We have deployed the constructed UrbanKGent-13B as a prototype system to offer urban knowledge graph construction service. The project service could be found in Github [^11] https://htmlpreview.github.io/?https://raw.githubusercontent.com/usail-hkust/UrbanKGent/main/UrbanKGent%20Demo/index.html. Figure 12 shows the screenshot of our demo system. Specifically, for the user, the system provides UrbanKGC service including relational triplet extraction and knowledge graph completion task. Following the instruction template proposed in this paper, the user can input urban text in the web shown in Figure 12 and the web will return extracted triplet. The user can also input two entities name and their geometry value, and the web serive will return completed geospatial relationship between them.
Appendix G Evaluation on Larger Dataset
As shown in Table 6, we derive UrbanKGent-13B for urban knowledge graph construction using constructed large-scale dataset in NYC and CHI. Specifically, we directly use the text record in NYC-Large and CHI-Large for the relational triplet extraction task. Then, we randomly sample the head-tail entity pairs (both of head and tail entities contain geometry information) from these triplets for knowledge graph completion. Nevertheless, iterating all head-tail entity pairs is a time-consuming task, so we just construct a KGC dataset consistent with the sacle of the RTE dataset. By performing RTE and KGC tasks, we obtain two large-scale UrbanKGs shown in Table 6. Compared with existing dataset and benchmark UUKG (Ning et al., 2023), we can clearly observe that our agent can only use one-fifth data to construct the UrbanKG with the same scale entities and triplets, but extend the relationship types to a thousand times. It is worth noting that all construction process is completed by a LLM agent without any mannul effort. We think it is the core of this work.
Figure 12. The prototype system of UrbanKGent.
Table 11. The experimental results of relational triplet extraction (RTE) and knowledge graph completion (KGC) on the NYC-Large dataset. To save the cost, We choose the best approaches in zero-shot-learning (ZSL), In-context-learning (ICL), Vanilla Fine-tuning (VFT), and UrbanKGent Inference setting.
Models GPT(acc/confidence) Human(acc)
RTE KGC RTE KGC
ZSL 0.36/4.01 0.45/3.95 0.42 0. 31
ICL 0.39/4.36 0.48/4.15 0.48 0.39
VFT 0.37/4.05 0.46/3.98 0.45 0.35
UrbanKGent
Inference
0.43/4.13 0.51/3.87 0.51 0.43
UrbanKGent-7B 0.44/4.27 0.50/4.07 0.53 0.44
UrbanKGent-13B 0.46/4.56 0.52/3.67 0.55 0.46
Appendix H Complexity and Effiency Analysis
Moreover, we also provide comprehensive efficiency analysis to show the latency and cost of different models when completing UrbanKGC tasks. Specifically, we report the total inference time and cost [^12] We subscribe to the NVIDIA A800 computing resources and GPT service from HKUST(GZ). Following the standard price instruction of HKUST(GZ), we could calculate cost of GPT-based baselines or Llama-based baselines. But in order to protect privacy of the university, we do not provide detailed prices information here. of each method completing with 1,000 RTE and KGC tasks. For the cost of GPT-4 service, we first count the number of prompt token and completion token spent on 1,000 RTE and KGC tasks, and then calculate the cost based on billing standards. As for the cost of UrbanKGent-13B, we first count the GPU running time spent on 1,000 RTE and KGC tasks, and then calculate the cost based on the A800 charging standard provided by HKUST(GZ).

### v2

@@ E.2 RQ2: Ablation Studies
UrbanKGent-7B 0.38/4.17 0.42/3.98 0.37 0.34
UrbanKGent-7B 0.34/4.06 0.45/4.02 0.34 0.39
UrbanKGent-7B 0.45/4.32 0.40/3.97 0.45 0.23
UrbanKGent-7B 0.44/4.10 0.47/3.85 0.46 0.43
E.3 RQ3: Complexity and Effiency Analysis
Table 11: Comparison among LLM-based UrbanKGC methods in four ways.
@@ E.3 RQ3: Complexity and Effiency Analysis
Moreover, we also provide comprehensive efficiency analysis to show the latency and cost of different models when completing UrbanKGC tasks. Specifically, we report the total inference time and cost [^8] We subscribe to the NVIDIA A800 computing resources and GPT service from HKUST(GZ). Following the standard price instruction, we could calculate cost of GPT-based baselines or Llama-based baselines. of each method completing with 1,000 RTE and KGC tasks. For the cost of GPT-4 service, we first count the number of prompt token and completion token spent on 1,000 RTE and KGC tasks, and then calculate the cost based on billing standards. As for the cost of UrbanKGent-13B, we first count the GPU running time spent on 1,000 RTE and KGC tasks, and then calculate the cost based on the A800 charging standard.
In addition, as reported in table 12, we also provide detailed inference latency of UrbanKGent family when deriving them for constructing different-scale of UrbanKGs.
Table 12: The inference latency comparison of UrbanKGC using UrbanKGent family. We use two middle-size dataset (i.e., NYC and CHI) and two large-scale dataset (i.e., NYC-Large and CHI-Large) for UrbanKG construction.
Dataset Latency (minutes) Data Volume
UrbanKGent-7B UrbanKGent-8B UrbanKGent-13B
NYC 1.19 1.58 3.19 2,089
CHI 0.62 1.13 1.68 1,102
NYC-Large 23.07 30.76 61.93 40,480
CHI-Large 16.55 21.93 44.47 28,868
- 0.57 0.76 1.53 Per 1,000 records
E.4 RQ4: Case Study
@@
E.5 RQ5: Agent Application
We have released the UrbanKGent family consisting of 7B, 8B and 13B version in the Huggingface. The opensourced UrbanKGent family offer urban knowledge graph construction service for the researcher in this field. We provide application example in New York and Chicago. Specifically, following the UrbanKGent Inference framework, we sequentially derive UrbanKGent-13B for relational triplet extraction and knowledge graph completion. The obtained initial UrbanKGs encodes diverse urban spatial, temporal and functional knowledge. Then, we propose to use a two-stage triplet filtering and relation merging operation to further improve the quality of constructed UrbanKGs.
In the first stage, low-frequency relations (occur 5 times or less) are merged into high-frequency relations based on relation similarity threshold. The remaining low-frequency triples, whose similarity with any high-frequency relation is below the threshold will be filtered out. In the stage two, we first perform relation clustering based on the embedding of relations. Then, Within each cluster, we prompt LLM to identify semantically similar relations that can be merged into a single relation category, resulting the final set of merged relations.
As shown in Table 13, both urban entity and relation can be pre-categorized into 4 coarse-grained ontologies: spatial, temporal, functional, and others. The entity percentage of spatial, temporal, functional and others in NYC-Large is (68.34%, 17.66%, 10.39%, 3.61%), and in CHI-Large is (63.29%, 16.68%, 12.09%, 7.94%). The relation percentage of spatial, temporal, functional and others in NYC-Large and CHI-Large are (57.36%, 16.38%, 20.89%, 5.37%) and (60.57%, 15.11%, 21.67% and 2.65%), respectively. After multi-view entity recognition and relation extraction (shown in Figure 4(a) in our paper), the fine-grained entity ontologies (6,281 and 2,559 entity types of NYC-Large and CHI-Large UrbanKGs, respectively) and fine-grained relation ontologies (2,138 and 1,366 relation types of NYC-Large and CHI-Large UrbanKGs, respectively) are obtained.
Table 13: The statistic of entity and relation ontology of constructed UrbanKGs on NYC-Large and CHI-Large datset.
UrbanKG Dataset
# Coarse-grained
Entity Ontology
# Fine-grained
Entity Ontology
# Coarse-grained
Relation Ontology
# Fine-grained
Relation Ontology
# Entity # Triplet
NYC-Large 4 6,281 4 2,138 228,928 905,442
CHI-Large 4 2,559 4 1,336 95,813 563,290
Appendix F Limitation and Future Work
This work has limitation on the further application demonstration of construction UrbanKGs, although proposed UrbanKGent family could construct a UrbanKG with a hundreds relationship using only one-fifth of data. In addition, the evaluation method in this work is cost-intensive although GPT evaluation and Human evaluation has been experimentally demonstrated to be consistent. Despite the above limitations, we hope the opensource UrbanKGC agent can foster more extensive UrbanKG research and broad smart city application. In the future, we will derive extra image-modality data to further enrich UrbanKGC.